A contractor has to supply 10,000 bearings per day to an automobile manufacturer . He finds that when he starts production run, he can produce 25,000 bearings per day. The cost of holding a bearing in stock for one year is Rs. 2 and the set up cost of a production run is Rs. 180 . Find the EOQ . How frequently should the production run he made ?


1
Expert's answer
2021-05-07T09:19:56-0400

This is the standard batch production model with "d=10,000, r=25,000,h=\dfrac{3}{365} \text{ and } s=1800"


"EOQ^*=\sqrt{\dfrac{2sd}{h[1-\frac{d}{r}]}}=\sqrt{\dfrac{2\times 1800\times \frac{2}{365}}{\frac{2}{365}[1-\frac{10,000}{25000}]}}=104,642"



"T^* =\dfrac{Q^*}{d} = 10.46" is the time between production runs.


A practical answer if "T^∗ = 10 \text{ with } EOQ^* = 100, 000."